# suppSupplementary References

### v1

@@ 3.1.3 Robustness Evaluation for Object Detection
To evaluate the robustness of VSSMs, transformers, and CNNs for object detection, we use COCO-O [8], COCO-DC [34], and COCO-C datasets. COCO-O contains 6,782 images and 26,624 labeled bounding boxes across six challenging domains: sketch, cartoon, painting, weather, handmake, and tattoo. COCO-DC is a subset of 1,127 images from the COCO 2017 validation set with diverse background changes introduced by a diffusion model. COCO-C comprises the COCO-2017 evaluation set, corrupted with common ImageNet-C style corruptions at various intensity levels. We assess performance using Average Precision (AP) and AP by object sizes (APs, APm, APl) on five models: ConvNext-T, Swin-T, Swin-S, VSSM-T, and VSSM-S.
@@
Figure 3: Mean Average Precision (mAP) score for different architectures on COCO-C dataset.
Model AP APs APm APl AP APs APm APl AP APs APm APl AP APs APm APl
               Original Color Texture Average
ConvNext-T 66.2 41.0 61.3 71.4 55.0 26.3 49.0 60.6 53.5 26.5 47.5 60.5 54.25 (-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.95)}}}$ 26.40(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-14.60)}}}$ 48.25(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-13.05)}}}$ 60.55(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.85)}}}$
Swin-T 66.3 45.1 61.6 71.6 55.1 29.2 48.0 62.0 53.6 31.4 45.2 61.4 54.35(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.95)}}}$ 30.30(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-14.80)}}}$ 46.60 (-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-15.00)}}}$ 61.70(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-09.90)}}}$
Swin-S 69.1 43.7 62.3 75.9 57.4 30.1 49.1 64.3 56.0 26.8 45.4 64.3 56.70(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-12.40)}}}$ 28.45(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-15.25)}}}$ 47.25(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-15.05)}}}$ 64.30(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.60)}}}$
VSSM-T 69.3 47.1 64.4 75.5 56.2 30.9 48.4 63.7 53.2 29.2 44.9 61.0 54.57(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-14.73)}}}$ 30.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-17.05)}}}$ 46.65(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-17.75)}}}$ 62.35(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-13.15)}}}$
VSSM-S 70.9 51.7 64.6 77.1 56.7 31.3 48.9 64.3 55.7 26.8 47.4 64.4 56.20(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-14.70)}}}$ 29.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-22.65)}}}$ 48.15(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-16.45)}}}$ 64.35(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-12.75)}}}$
Figure 4: Mean Intersection over Union (mIoU) score for different architectures on AED20k-C dataset.
@@
3.1.4 Robustness Evaluation for Semantic Segmentation
To assess the robustness of segmentation models, we employ the ADE20K dataset [48]. We corrupt 2,000 images from the ADE20K validation set using ImageNet-C [17] corruptions at various intensity levels. The performance of the models is evaluated using the mean Intersection over Union (mIoU).
We observe a similar trend in the performance of segmentation models on the corrupted ADE20K dataset as seen with object detection models on COCO-C. Fig. 4 shows that VSSM-T and VSSM-S consistently outperform their Swin counterparts across various ImageNet-C corruptions in object segmentation, demonstrating superior robustness. The high performance of VSSM models on the original dataset also transfers effectively to the corrupted version of the dataset. Notably, even the smallest VSSM-T model surpasses the larger Swin-S model in most of the corruption scenarios.
@@ 3.2 Robustness against Adversarial Attacks
Table 7: Robust accuracy of models under white-box and black-box settings for FGSM and PGD attacks. Adversarial examples crafted on surrogate models are used to evaluate robustness of target models.
@@
VSSM-T
VSSM-S
VSSM-B
@@
VSSM-T 42.90 66.34 65.10 80.20 72.66 73.80 72.84 79.46 83.64 86.94 72.22 76.38 74.60
VSSM-S 62.24 48.42 63.00 79.70 71.40 71.04 70.18 78.42 81.56 84.58 70.96 72.92 71.14
VSSM-B 65.52 66.96 51.24 81.22 73.54 73.20 72.32 79.12 83.62 86.24 73.24 76.30 73.88
ResNet-50 81.38 83.24 83.84 30.46 80.30 82.20 83.38 75.94 84.74 89.12 80.64 85.00 85.42
ConvNext-T 69.00 71.46 71.18 77.96 36.36 61.96 63.76 76.92 82.74 85.58 67.06 71.88 70.78
ConvNext-S 69.54 70.62 70.48 79.48 63.48 49.10 63.62 78.48 82.94 85.02 69.24 71.54 69.74
ConvNext-B 70.54 71.78 69.78 81.72 67.34 66.24 51.32 80.26 83.98 86.22 69.86 73.44 70.84
ViT-T 85.92 88.30 88.88 82.68 85.08 86.56 88.06 2.28 50.04 69.90 75.66 79.72 82.32
ViT-S 81.74 83.76 84.46 81.70 82.34 82.78 84.38 45.40 11.02 54.82 72.50 75.98 77.88
ViT-B 82.46 84.02 84.48 81.90 82.04 82.24 83.62 58.50 53.84 24.24 75.82 77.62 78.54
Swin-T 72.40 76.06 75.58 82.82 71.46 72.12 71.96 76.48 80.66 85.46 28.86 56.22 55.12
Swin-S 78.24 79.10 79.30 85.46 77.96 77.10 77.02 79.38 83.04 86.58 61.94 48.00 63.90
Swin-B 78.88 79.40 79.04 86.60 78.16 78.46 77.32 81.48 84.74 87.64 66.82 68.28 54.76
Projected Gradient Decent (PGD) at $\epsilon=\frac{8}{255}$
VSSM-T 0.00 37.26 35.56 81.44 63.54 67.18 65.88 82.58 89.16 91.80 65.76 74.72 71.50
VSSM-S 16.88 0.02 23.94 77.54 56.24 57.72 56.16 79.26 86.06 89.42 58.92 65.92 61.46
VSSM-B 15.98 25.26 0.00 77.38 56.30 57.94 53.42 79.76 86.24 89.94 57.40 66.52 61.74
ResNet-50 83.34 86.98 87.04 0.00 81.00 84.92 86.30 82.48 89.96 93.18 84.18 89.24 89.12
ConvNext-T 60.58 67.06 66.48 79.76 0.00 31.32 39.70 80.44 87.28 90.46 56.02 66.62 64.40
ConvNext-S 58.04 63.20 62.96 78.22 24.96 0.00 33.96 80.02 86.20 89.86 52.70 62.58 60.54
ConvNext-B 58.24 64.04 60.40 80.80 35.84 36.60 0.00 81.48 87.22 90.36 51.80 63.16 58.62
ViT-T 92.10 93.82 94.26 88.14 90.44 92.18 93.28 0.00 55.58 77.34 84.52 88.16 89.24
ViT-S 87.28 89.34 89.70 85.96 86.34 87.90 89.26 38.16 0.00 48.68 76.98 80.34 82.48
ViT-B 86.18 87.74 87.72 85.04 85.56 86.10 87.62 53.50 37.48 0.00 77.02 79.56 80.86
Swin-T 83.10 87.58 87.06 89.58 79.30 81.52 80.28 83.32 89.82 92.80 0.00 39.00 37.88
Swin-S 83.08 86.28 86.26 89.76 79.60 80.68 79.84 82.96 88.80 92.38 25.40 0.00 31.44
Swin-B 82.52 85.12 84.78 89.88 79.18 80.08 78.92 84.24 89.74 92.86 27.30 34.36 0.02
Filter $\rightarrow$ Low-Pass High-Pass All-Pass
VSSM-T 96.46 96.04 95.22 93.02 91.70 91.18 58.12 32.50 16.46 5.74 2.92 2.82 24.00 5.16 0.28
VSSM-S 97.02 96.90 96.48 94.10 92.62 92.24 67.64 44.00 28.18 11.36 6.46 5.48 32.54 10.90 0.84
VSSM-B 97.28 97.00 96.56 94.28 92.90 92.24 66.60 45.44 27.68 10.36 6.08 4.78 33.10 9.88 0.40
@@
Adversarial Attacks in Spatial Domain: We conduct experiments in both white-box and black-box settings using the Fast Gradient Sign Method (FGSM) [13] and Projected Gradient Descent (PGD) [32] with an $l_{\infty}$-norm and $\epsilon=8/255$. FGSM is a single-step process, while PGD operates as a multistep method, iterating for 20 steps with a step size of $2/255$. Tab. 7 (top) displays the robust accuracy scores under white-box (diagonal entries) and black-box (off-diagonal entries) adversarial attacks for FGSM. VSSM-T and VSSM-S exhibit higher white-box attack robustness compared to their Swin Transformer counterparts, but this pattern does not extend to the larger VSSM-B model. This indicates that VSSM’s robustness advantage over Swin Transformers may not consistently scale with increased model size. For black-box settings, generally attacks transfer more easily within the same architecture family (e.g., from VSSM-T to VSSM-S) than across different architectures (e.g., from VSSM to Swin Transformer). With stronger PGD-20 attack (Tab. 7 (bottom)), we observe all the models performance drops to zero under white-box setting, revealing their vulnerability to stronger adversarial attacks. Transferability in black-box setting for PGD attack shows similar trend to that obtained for FGSM attack. For further analysis across other model families, refer to Appendix A.
@@
Adversarial Fine-tuning on Downstream Datasets: We adversarially fine-tune ImageNet pretrained VSSM, ViT, and Swin on downstream datasets using the TRADES [47] objective with varying robustness strength $\beta$ and an $l_{\infty}$ perturbation budget $\epsilon=8$. In Fig. 5, we plot the clean and robust accuracy under PGD-100 at $\epsilon=8$. On Imagennette [1], VSSM-T shows strong performance in both clean and robust accuracy across different $\beta$ levels, followed closely by Swin-T, while ViT-T exhibits the lowest robustness. However, on the low-resolution CIFAR-10 dataset with significantly lower number of patches, attention-based ViT models perform better than Mamba-based VSSM models.
@@ 4 Conclusion
We offer further insights into our study through various sections in the appendix, incorporating both versions of the VSSM models[28], v2 (reported in the main paper) and v0. Firstly, we expand our analysis on adversarial attacks across recent CNN and Transformer models in Section A, presenting results for white-box and transfer-based black-box attacks. In Section B, we provide detailed results on non-adversarial corruptions. Section B.1 explores information drop along the scanning direction. Sections B.2 and B.4 offer additional detailed results on Salient and Non-Salient Patch Drop and Random Patch Drop, respectively. Section B.3 presents further results on patch shuffling. Additionally, Section B.5 provides results on ImageNet-E, ImageNet-B, ImageNet-V2, ImageNet-A, ImageNet-R, and ImageNet-S. In Section B.6, we report results for each common corruption method across different models.
Appendix A Adversarial Transferability

### v2

@@ 3.1.3 Robustness on Object Detection
We evaluate VSSM, transformer, and CNN robustness for object detection using COCO-O [8], COCO-DC [35], and COCO-C datasets. COCO-O includes 6,782 images with 26,624 bounding boxes across six domains (sketch, cartoon, painting, weather, handmake, tattoo). COCO-DC contains 1,127 COCO 2017 validation images with diffusion model-induced background changes. COCO-C applies ImageNet-C style corruptions to the COCO-2017 evaluation set at various intensities. Performance is assessed using Average Precision (AP) and AP by object sizes (APs, APm, APl) on five models: ConvNext-T, Swin-T, Swin-S, VMamba-T, and VMamba-S.
@@
Figure 3: Performance comparison of different architectures on the AED20k-C and COCO-C datasets for segmentation (top) and detection (bottom) tasks. The top figure shows the Mean Intersection over Union (mIoU) score on the AED20k-C dataset, while the Mean Average Precision (mAP) score on the COCO-C dataset.
@@
3.1.4 Robustness on Semantic Segmentation
We assess segmentation model robustness using 2,000 images from the ADE20K [49] validation set, corrupted with ImageNet-C [19] at various intensities. Performance is evaluated using mean Intersection over Union (mIoU).
Fig. 3 shows VMamba-T and VMamba-S consistently outperforming Swin counterparts across various ImageNet-C corruptions in segmentation, mirroring trends seen in object detection. The high performance of VMamba models on the original dataset also transfers effectively to the corrupted version of the dataset. Notably, VMamba-T surpasses the larger Swin-S in most corruption scenarios. Additional results are provided in Appendix B.4.
@@ 3.2 Robustness against Adversarial Attacks
Table 5: Robust accuracy of models under white-box and black-box settings for FGSM attack. Adversarial examples crafted on surrogate models are used to evaluate robustness of target models.
@@
VMamba-T
VMamba-S
VMamba-B
MambaVision-T
MambaVision-S
MambaVision-B
@@
VMamba-T 42.90 66.34 65.10 74.44 73.00 73.16 80.20 72.66 73.80 72.84 79.46 83.64 86.94 72.22 76.38 74.60
VMamba-S 62.24 48.42 63.00 71.84 71.92 71.34 79.70 71.40 71.04 70.18 78.42 81.56 84.58 70.96 72.92 71.14
VMamba-B 65.52 66.96 51.24 75.22 73.82 73.08 81.22 73.54 73.20 72.32 79.12 83.62 86.24 73.24 76.30 73.88
MambaVision-T 77.74 79.04 79.78 46.18 73.30 75.64 82.06 79.50 80.88 81.82 78.30 84.20 87.86 77.78 81.24 80.80
MambaVision-S 77.86 79.56 79.24 74.96 53.42 73.84 83.60 80.40 81.14 81.26 79.90 85.64 88.52 78.50 82.40 81.88
MambaVision-B 75.90 77.84 77.00 75.46 71.96 52.68 83.46 80.24 80.52 79.92 79.82 85.18 88.12 78.46 81.60 80.30
ResNet-50 81.38 83.24 83.84 81.06 82.54 84.88 30.46 80.30 82.20 83.38 75.94 84.74 89.12 80.64 85.00 85.42
ConvNext-T 69.00 71.46 71.18 73.50 73.92 74.40 77.96 36.36 61.96 63.76 76.92 82.74 85.58 67.06 71.88 70.78
ConvNext-S 69.54 70.62 70.48 74.68 74.38 75.24 79.48 63.48 49.10 63.62 78.48 82.94 85.02 69.24 71.54 69.74
ConvNext-B 70.54 71.78 69.78 76.78 75.02 74.82 81.72 67.34 66.24 51.32 80.26 83.98 86.22 69.86 73.44 70.84
ViT-T 85.92 88.30 88.88 82.16 85.06 87.24 82.68 85.08 86.56 88.06 2.28 50.04 69.90 75.66 79.72 82.32
ViT-S 81.74 83.76 84.46 78.94 80.78 82.68 81.70 82.34 82.78 84.38 45.40 11.02 54.82 72.50 75.98 77.88
ViT-B 82.46 84.02 84.48 80.40 81.82 82.46 81.90 82.04 82.24 83.62 58.50 53.84 24.24 75.82 77.62 78.54
Swin-T 72.40 76.06 75.58 76.80 77.16 78.06 82.82 71.46 72.12 71.96 76.48 80.66 85.46 28.86 56.22 55.12
Swin-S 78.24 79.10 79.30 80.12 81.90 81.40 85.46 77.96 77.10 77.02 79.38 83.04 86.58 61.94 48.00 63.90
Swin-B 78.88 79.40 79.04 81.48 82.16 81.04 86.60 78.16 78.46 77.32 81.48 84.74 87.64 66.82 68.28 54.76
@@
Adversarial Attacks in Spatial Domain: We conduct experiments in both white-box and black-box settings using the Fast Gradient Sign Method (FGSM) [13] and Projected Gradient Descent (PGD) [33] with an $l_{\infty}$-norm and $\epsilon=8/255$. FGSM is a single-step process, while PGD operates as a multistep method, iterating for 20 steps with a step size of $2/255$. Tab. 5 (top) displays the robust accuracy scores under white-box (diagonal entries) and black-box (off-diagonal entries) adversarial attacks for FGSM. ‘T’ and ‘S’ versions of VMamba and MambaVision models exhibit higher white-box attack robustness compared to their Swin Transformer counterparts, but this pattern does not extend to the larger ‘B’ models. This indicates that VSSM’s robustness advantage over Swin Transformers may not consistently scale with increased model size. For black-box settings, attacks transfer rate is high within the same architecture family than across different architectures. As expected, under stronger iteraive attack (PGD), all models’ performance almost drops to zero in white-box settings. For the black-box transferability, we observe similar trends to FGSM attack (see Appendix C).
@@
Adversarial Fine-tuning on Downstream Datasets: We adversarially fine-tune ImageNet pretrained VSSM, ViT, and Swin on downstream datasets using the TRADES [48] objective with varying robustness strength $\beta$ and an $l_{\infty}$ perturbation budget $\epsilon=\frac{8}{255}$. In Fig. 4, we plot the clean and robust accuracy under PGD-100 at $\epsilon=\frac{8}{255}$. On Imagennette [1], VSSM-T shows strong performance in both clean and robust accuracy across different $\beta$ levels, followed closely by Swin-T, while ViT-T exhibits the lowest robustness. However, on the low-resolution CIFAR-10 dataset with significantly lower number of patches, ViT models perform better than Mamba-based VSSM models.
@@ 4 Conclusion
We provide additional insights into our study through various sections in the appendix, which include results for both versions of the VMamba models [29]: v2 (reported in the main paper) and v0. Results that do not specify the VMamba version should be assumed to refer to v2 by default.
@@
In Section A, we provide detailed results on Information Drop experiments. Section A.1 explores information drop along the scanning direction. Sections A.2 and A.3 offer additional detailed results on Salient and Non-Salient Patch Drop and Random Patch Drop, respectively. Section A.4 presents further results on patch shuffling. Section B provides results on Image Corruptions. In Section B.1, robustness against global corruptions such as common corruption, and out-of-distribution datasets is evaluated. Additionally, in Section B.2, we provide model calibration results on in-distribition and out-of-distribution datasets. In Section B.3, we report results on fine-grained corruptions across different models. In Section B.4, we report image corruption results for the task of object detection and semantic segmentation. Finally, we expand our analysis of Adversarial Attacks across models in Section C. We present results for white-box attacks, transfer-based black-box attacks, and frequency-based attacks.
Appendix A Robustness against Information Drop
A.1 Information Drop along the Scanning Axis
A.2 Random Patch Drop
In Table 6 we expand the random patch drop experiments from the main paper (Table 1) to patch sizes 56$\times$56 and 224$\times$224. Furthermore, in Figure 9, 10, 11, 12, 13, and 14 we expand our analysis on random patch drop to several other models.
A.3 Salient and Non-Salient Patch Drop
In Figure 15 and 16, we report results on salient and non-salient patch drop of information.
A.4 Patch Shuffling
Appendix B Robustness against Image Corruptions
B.1 Robustness against Global Corruptions
In Table 7 relative corruption error is reported across models discussed in the main paper. In Figure 18 and 19, we present the relative corruption error and mean corruption error (mCE) of various models subjected to all 19 corruption methods applied to the ImageNet dataset. The common corruption consists of various types of synthetic corruptions to assess the models’ robustness against image distortions like noise, blur, and compression artifacts. In Table 8 we report results on ImageNetV2 [42], ImageNet-A [20], ImageNet-R [18], and ImageNet-S [11]. These datasets are designed to evaluate the robustness and generalization capabilities of models trained on the original ImageNet dataset. ImageNetV2 [42] tests the models’ performance on previously unseen images that follow the same distribution as the training data, while ImageNet-A [20] contains naturally occurring adversarial examples that are difficult for models to classify correctly. ImageNet-R [18] consists of images with different artistic renditions to test the models’ ability to generalize to different visual domains and styles, and ImageNet-S [11] consists of sketch-based images.
B.2 Model Calibration
Model calibration assesses how well a model’s predicted confidence aligns with its actual accuracy. For example, if a model predicts a confidence level of 70% for its predictions, a well-calibrated model should have an actual accuracy close to 70%. To quantify this alignment, we use the Expected Calibration Error (ECE). ECE involves dividing predictions into $M$ bins based on their confidence levels (e.g., 60%-70%, 70%-80%). For each bin, the average accuracy and confidence are computed, and the ECE is the weighted average of the differences between these values. Calibration can also be evaluated visually using reliability diagrams, which plot predicted confidence against actual accuracy; a well-calibrated model should show points near the diagonal. Additionally, confidence histograms reveal the distribution of prediction confidences. Evaluation is performed on both in-distribution data (e.g., ImageNet, ImageNetV2) and out-of-distribution data (e.g., ImageNet-R, ImageNet-S, ImageNet-A), with $M=15$ bins used in all experiments. In Figure 20, 21, and 22, we report the ECE plot the reliability diagrams for Tiny, Small, and Base version of different architectures, respectively. Figure 23, we report the ECE score. We observe that while ViT models report a low ECE error on in-distribution datasets, the error increases significantly for out-of-distribution datasets. An exception is the Base model, which can be attributed to the ViT-B model being pretrained on a larger dataset (ImageNet-21k) comapared to other models. For the hybrid-based VSSM model MambaVision we observe a contrasting behaviour with ViT models, while the ECE score on in-distribution datasets is relatively high, it improves significantly on out-of-distribution datasets. The improvement in ECE for MambaVision on out-of-distribution datasets suggests that the hybrid approach of combining vision transformers with other model components may enhance calibration performance in these scenarios.
B.3 Robustness against Fine-grained Corruptions
In Table 9 (left) we report results on ImageNet-E dataset across several CNNs, Transformers, and VSSM-based models, and in Table 9 (right), we report results on ImageNet-B dataset.
B.4 Robustness Evaluation for Object Detection and Semantic Segmentation
In Table 10, we report AP scores of different architectures on COCO-DC dataset. On COCO-DC, we observe that VMamba models’ high performance on clean images does not translate to color and texture background variations. Swin-S model achieves the highest average AP score of $56.70$ across the background variations, followed by score of $56.20$ by VMamba-S model. Figure 24 shows the mIoU, mAcc, and aAcc scores on ADE-20K after applying common corruptions to the dataset. Similarly, Figure 25 displays the mAP, APs, APm, and APl scores on the COCO-C dataset. All the scores are averaged across all severity levels of each corruption.
Appendix C Robustness against Adversarial Attacks